Grade 12Mechanics

A man teavels 30m north and 20m towards east and finally travels 30^2m towards southwest find the displacement and direction from the origin.

wani
9 Years agoGrade 12

1 Answer

Naman singh
9 Years ago
Draw the figure and apply Pythagoras theorem900^2=20^2+x^2x=899.7777Now subtract 30m from xAns 869.77